Emergency!!!!! Buried and need help in understanding these problems.
(1) Find the x-intercept and y-intercept of the equation 5x -3y = 30.
(2) Find the equation for the line with slope .5 and y-intercept 3.
(3) Find the equation for the line with y-intercept 3 that is perpindicular to the line y = 2/3 x -4. I came up with 3y = 9 -2x am I close on this one?
(4) If f(x) =x^2 + 5, find f(a+h) - f(a). My solution was 2ah+h^2+10 is that correct?

Answer by fractalier(6550)

Please list them one at a time next time...tutors will help you more readily...
1) x = 6, y = -10
2) y = (1/2)x + 3
3) y = (3/2)x + 3
4) f(a+h) - f(a) = (a+h)^2 + 5 - (a^2 + 5) = 2ah + h^2